# Enhancing Robustness in Deep Reinforcement Learning: A Lyapunov Exponent Approach

### v1

@@ 4 Chaotic rewards
In this section, we show that chaotic state trajectories can also impact a policy’s performance and produce chaotic reward trajectories as determined by the Maximal Lyapunov Exponent ($\lambda_{1}$). This instability creates a fractal return surface in which small state perturbations produce significantly different total rewards. We argue that adversarial attack methods could leverage these high-frequency oscillations, repeatedly injecting perturbations which cause the agent to follow the state trajectories that attain the lowest total reward. This lack of robustness poses a significant problem for real-world control systems where the worst-case performance is often more significant than the average performance.
By considering the reward over a state trajectory as a trajectory in a one-dimensional reward space, the stability of the reward can also be measured using Lyapunov Exponents. Given this space is one-dimensional only one Lyapunov Exponent exists; however, this single exponent can still be used to reliably identify the stability of reward trajectories. Negative $\lambda_{1}$ values indicate that small perturbations to the system’s state still produce converging long-term rewards. Moreover, for a bounded reward function, the reward trajectories cannot diverge indefinitely; thus a positive $\lambda_{1}$ indicates long-term reward is chaotic as small changes to the system state produce exponentially diverging bounded reward trajectories.
Figure 5: Reward MLE interquartile mean for the Pointmass (PM), Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (SW), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) when controlled by SAC, TD3 and Dreamer V3 (DR3). Each policy-environment combination is independently trained with three random seeds and the reward MLE for each seed is calculated using 20 initial states. A bootstrapped 95% confidence interval is included to show the variation in reward stability across random seeds.
@@
Figure 5 provides the average reward MLE and bootstrapped 95% confidence interval for each policy-environment pair defined in Section 3. These graphs indicate that the reward trajectories produced in simple low-dimensional environments (Pointmass, Cartpole Balance, Cartpole Swingup) are stable as they have negative reward MLE. As a result, a single small state perturbation does not harm the total reward attained as the subsequent reward trajectories converge. Conversely, in high-dimensional control systems (Walker Stand, Walker Walk, Walker Run and Cheetah Run), the reward MLE is positive indicating the reward is highly sensitive to arbitrarily small changes in the system state. Consequently, reward trajectories in these systems are unstable as similar initial states produce significantly different sequences of rewards. Plotting the reward trajectories for one of these chaotic interactions (Figure 6) shows that a single perturbation with varying magnitude can have a huge detrimental impact on the total reward attained despite performing well on average. This poses a significant problem for real-world applications of RL as it is impossible to guarantee worst-case performance in a chaotic control system.
@@ 5 Maximal Lyapunov Exponent regularisation
Dynamcis Predictor $(q_{\phi}:\mathcal{H}\rightarrow\mathcal{P}(\mathcal{Z}))$
@@
$\mathcal{L}^{\lambda_{1}}(\theta)\leftarrow 0$eqparboxCOMMENT# Initlaise the MLE regularisation loss
@@
We base our regularisation on Dreamer V3 [9], a general-purpose model-based RL algorithm which attains state-of-the-art performance across a diverse set of control tasks. To achieve this, Dreamer V3 uses a Recurrent State Space Model (RSSM) consisting of an Encoder $(q_{\phi}:\mathcal{S}\times\mathcal{H}\rightarrow\mathcal{P}(\mathcal{Z}))$, Decoder $(p_{\phi}:\mathcal{H}\times\mathcal{Z}\rightarrow\mathcal{P}(\mathcal{S}))$, Dynamics Predictor $(q_{\phi}:\mathcal{H}\rightarrow\mathcal{P}(\mathcal{Z}))$ and Sequence Model $(f_{\phi}:\mathcal{H}\times\mathcal{Z}\times\mathcal{A}\rightarrow\mathcal{H})$ to predict state trajectories ($s_{t}$), bootstrapped $\lambda$-return trajectories ($R_{t}^{\lambda}$) [25] and state value trajectories ($v_{\phi}(s_{t})$) over a short time horizon $T$. The policy is then trained to maximise the normalised advantage estimates using REINFORCE gradients [30] and an entropy regulariser ($\text{H}[\cdot]$[29] with weighting coefficient $\eta$. The full loss function used to train Dreamer V3’s policy is outlined in Equation 4.
$\mathcal{L}^{\text{Dr3}}(\theta)\doteq-\sum_{t=1}^{T}\text{sg}\left(\frac{R_{t}^{\lambda}-v_{\phi}(s_{t})}{\text{max}(1,S)}\right)\log\pi_{\theta}(a_{t}|s_{t})+\eta\text{H}\left[\pi_{\theta}(a_{t}|s_{t})\right]$ (4)
$\mathcal{L}^{\lambda_{1}}(\theta)\doteq\sum_{t=1}^{T}\left(\mathop{\text{Var}}_{L}(S_{t})+\mathop{\text{Var}}_{L}(H_{t})\right)$ (5)
@@
At its core, Dreamer V3 uses a stochastic RSSM to predict the state and reward trajectories over a predefined time horizon given an initial starting state $s_{0}$ and an internal representation $h_{0}$. Due to the stochastic nature of this model, repeating the same trajectory predictions $L\in\mathbb{N}$ times produces a set of state trajectories ($S_{t}=\left\langle s_{t,1},s_{t,2},...,s_{t,L}\right\rangle$) and internal representation trajectories ($H_{t}=\left\langle h_{t,1},h_{t,2},...,h_{t,L}\right\rangle$), each of which provides a plausible estimate of the future states. The variance between trajectories ($\mathop{\text{Var}}_{L}(\cdot)$) thus provides an estimation of the local state divergence as the state perturbation size approaches 0. Therefore, to minimise $\lambda_{1}$ and improve the stability of Dreamer V3 subject to state perturbation, we propose incorporating the regularisation term outlined in Equation 5 into the policy loss (Equation 6). Including this regularisation term as an additional weighted term forces agents to consider the stability of the system during the optimisation process. This incentivises the policy to produce stable state trajectories which attain high rewards instead of solely optimising the expected return. The complete algorithm for calculating the MLE regularisation term is provided in Algorithm 1 using the notation outlined by Hafner et al [9].
@@ 6 Experiments
In this section, we investigate the impact of the proposed MLE regularisation has on Dreamer V3. We show that the inclusion of this term reduces the chaotic state dynamics produced by the control policy and that this improved stability increases performance when noise is introduced. For these experiments, we train three instances of Dreamer V3 with MLE regularisation and reuse the SAC, TD3 and Dreamer V3 policies from sections 3&4. When estimating state divergence, the RSSM predicts $L=3$ plausible future trajectories over which the state and internal representation variance is measured. Increasing $L$ will produce a more accurate estimate of state divergence; however, this will require more computational resources. Therefore, to maintain a similar training time to that of Dreamer V3 we set $L=3$. All other hyperparameters are consistent with the Dreamer V3 baseline.
@@
Table 2: Average Total reward and average MLE produced when Dreamer V3 (DR3) and Dreamer V3 with MLE regularisation (MLE DR3) control various environments sampled from the DeepMind Control Suite [27]. Each policy-environment combination is independently trained with three random seeds using the hyperparameters outlined in Appendix A.2.
@@
Figure 7: Total episode reward for the Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (SW), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) environments when controlled by trained instances of SAC, TD3, Dreamer V3 (DR3) and Dreamer V3 with MLE regularisation (MLE DR3) subject to $\mathcal{N}(0,\sigma)$ Gaussian observation noise. Each policy-environment combination is independently trained with three random seeds and the average episode reward with a bootstraped 95% confidence interval is reported over 80 evaluation episodes each with a fixed length of 1000 steps.
To identify how robust these regularised Dreamer V3 policies are to continual state perturbations we measure their performance when Gaussian noise is added to each observation. Each policy is trained without observation noise ($\sigma=0$) in the fully deterministic variant of each environment and then tested with Gaussian noise ($\mu=0$ and $\sigma\in[0,1]$). Performance is measured over 240 episodes with a fixed length of 1,000 steps and a maximum reward per step of 1. Figure 7 shows the interquartile mean and bootstrapped 95% confidence interval for the total reward attained by each policy-environment interaction subject to various levels of Gaussian noise. This shows MLE regularisation significantly improves the performance of Dreamer V3 in four of the noisy control systems while the other two environments (Walker Stand and Cheetah Run) attain similar performance as the Dreamer V3 baseline. Furthermore, examining the state trajectories produced by Dreamer V3 and Dreamer V3 + MLE regularisation when controlling the Walker Stand task (Figure 8) shows the regularisation improves the stability of the control interaction as the trajectories do not diverge significantly. These findings indicate MLE regularisation can improve the robustness of Dreamer V3 subject to observation noise as it produces more consistent and reliable behaviour in the face of uncertainties.
@@ 7 Conclusion
A key issue preventing the application of deep Reinforcement Learning to real-world environments is the need for guaranteed stability and performance in the face of noisy observations and adversarial attacks. In this work, we set out to identify the impact a single perturbation has on the long-term behaviour of deep RL policies in continuous control environments. Using the spectrum of Lyapunov Exponents we established the MDP can produce chaotic state and reward trajectories which are highly sensitive to initial conditions. This instability poses two threats to the application of deep RL to real-world problems where it is infeasible to attain an accurate measurement of the system state. First, small state perturbation can have a large impact on the performance of trained deep RL policies, even where the average performance is good. This can create hazards in real-world conditions where observation noise is prevalent and can be exploited by adversarial attack methods. Second, even when deep RL policies perform well, they can produce unpredictable behaviours which is undesirable in most real-world applications.
To mitigate these chaotic dynamics and improve robustness we propose Maximal Lyapunov Exponent regularisation. This novel approach uses the recurrent state space model to estimate the local state divergence and incorporates this into the Dreamer V3’s policy loss. In effect, the agent optimises its confidence in future trajectories jointly with its expectations of rewards. While MLE regularisation helps improve the robustness of the agent’s policies, this approach assumes an accurate estimation of the local state divergence. In environments where the RSSM struggles to capture state dynamics, the effectiveness of the proposed regularization may be diminished. However, in our experiments, we demonstrate that this regularisation improves the stability of the learnt policies, thereby making them more robust to state perturbations.
@@ Appendix A Appendix
A.1 State & action space composition
Table 3: State space definition for each environment used in this paper. Each dimension of the state space represents a unique aspect of the control system and can contain any real value. Positions are measured in meters ($m$), angles are measured in radians ($rad$), velocities are measured in meters per second ($m/s$) and angular velocities are measured in radians per second ($rad/s$).
@@ A.1 State & action space composition
x & y Position
x & y Velocity
@@ B.1 Default values
Table 5: Parameters used to calculate the spectrum of Lyapunov Exponents using the method outlined by Benettin et al [2, 3].
Parameter Default Value
@@ B.2 Number of iterations ablation study
Figure 9: Estimated Maximal Lyapunov Exponent of environments sampled from the DeepMind Control Suite when controlled by SAC, TD3 and Dreamer V3 (DR3). Each policy-environment combination is independently trained with three random seeds and evaluated using the parameters in Table 5 except the number of iterations, which varies from 1 to 100. The mean and 95% confideince interval indicate that MLE converges after 100 iterations.
@@ B.3 Number of samples ablation study
Figure 10: Estimated Maximal Lyapunov Exponent of environments sampled from the DeepMind Control Suite when controlled by SAC, TD3 and Dreamer V3 (DR3). Each policy-environment combination is independently trained with three random seeds and evaluated using the parameters in Table 5 except the number of initial samples, which varies from 1 to 20. The mean and 95% confideince interval indicate that MLE converges with 20 initial samples.

### v2

@@ 4 Chaotic rewards
In this section, we show that chaotic state trajectories can also impact a policy’s performance and produce chaotic reward trajectories as determined by the Maximal Lyapunov Exponent. This instability creates a fractal return surface in which small state perturbations produce significantly different total rewards. We argue that adversarial attack methods could leverage these high-frequency oscillations, repeatedly injecting perturbations which cause the agent to follow the state trajectories that attain the lowest total reward. This lack of robustness poses a significant problem for real-world control systems, where the worst-case performance is often more significant than the average performance.
By considering the reward over a state trajectory as a trajectory in a one-dimensional reward space, the stability of the reward can also be measured using Lyapunov Exponents. Given that this space is one-dimensional, only one Lyapunov Exponent exists; however, this single exponent can still be used to reliably identify the stability of reward trajectories. Negative $\lambda_{1}$ values indicate that small perturbations to the system’s state still produce converging long-term rewards. Moreover, for a bounded reward function, the reward trajectories cannot diverge indefinitely; thus a positive $\lambda_{1}$ indicates that the long-term reward is chaotic as small changes to the system state produce exponentially diverging bounded reward trajectories.
Figure 5: Reward MLE interquartile mean for the Pointmass (PM), Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (WS), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) when controlled by SAC, TD3 and Dreamer V3 (DR3). Each policy-environment combination is independently trained with three random seeds and the reward MLE for each seed is calculated using 20 initial states. A bootstrapped 95% confidence interval is included to show the variation in reward stability across random seeds.
@@
Figure 5 provides the average reward MLE and bootstrapped 95% confidence interval for each policy-environment pair defined in Section 3. These graphs indicate that the reward trajectories produced in simple low-dimensional environments (Pointmass, Cartpole Balance, Cartpole Swingup) are stable as they have negative reward MLE. As a result, a single small state perturbation does not harm the total reward attained as the subsequent reward trajectories converge. Conversely, in high-dimensional control systems (Walker Stand, Walker Walk, Walker Run and Cheetah Run), the reward MLE is positive, indicating that the reward is highly sensitive to arbitrarily small changes in the system state. Consequently, reward trajectories in these systems are unstable, as similar initial states produce significantly different sequences of rewards. Plotting the reward trajectories for one of these chaotic interactions (Figure 6) shows that a single perturbation with varying magnitude can have a huge detrimental impact on the total reward attained despite performing well on average. This poses a significant problem for real-world applications of RL as it is impossible to guarantee worst-case performance in a chaotic control system.
@@ 5 Maximal Lyapunov Exponent regularisation
Dynamics Predictor $(q_{\phi}:\mathcal{H}\rightarrow\mathcal{P}(\mathcal{Z}))$
@@
$\mathcal{L}^{\lambda_{1}}(\theta)\leftarrow 0$eqparboxCOMMENT# Initialise the MLE regularisation loss
@@
We base our regularisation on Dreamer V3 [9], a general-purpose model-based RL algorithm which attains state-of-the-art performance across a diverse set of control tasks. To achieve this, Dreamer V3 uses a Recurrent State Space Model (RSSM) consisting of an Encoder $(q_{\phi}:\mathcal{S}\times\mathcal{H}\rightarrow\mathcal{P}(\mathcal{Z}))$, Decoder $(p_{\phi}:\mathcal{H}\times\mathcal{Z}\rightarrow\mathcal{P}(\mathcal{S}))$, Dynamics Predictor $(q_{\phi}:\mathcal{H}\rightarrow\mathcal{P}(\mathcal{Z}))$ and Sequence Model $(f_{\phi}:~{}\mathcal{H}\times\mathcal{Z}\times\mathcal{A}\rightarrow\mathcal{H})$ to predict state trajectories ($s_{t}$), bootstrapped $\lambda$-return trajectories ($R_{t}^{\lambda}$) [25] and state value trajectories ($v_{\phi}(s_{t})$) over a short time horizon $T$. The policy is then trained to maximise the normalised advantage estimates using REINFORCE gradients [29] and an entropy regulariser ($\text{H}[\cdot]$[30] with weighting coefficient $\eta$. The full loss function used to train Dreamer V3’s policy is outlined in Equation 4.
$\mathcal{L}^{\text{Dr3}}(\theta)\doteq-\sum_{t=1}^{T}\left[\text{sg}\left(\frac{R_{t}^{\lambda}-v_{\phi}(s_{t})}{\text{max}(1,S)}\right)\log\pi_{\theta}(a_{t}|s_{t})+\eta\text{H}\left[\pi_{\theta}(a_{t}|s_{t})\right]\right]$ (4)
$\mathcal{L}^{\lambda_{1}}(\theta)\doteq\sum_{t=1}^{T}\left[\mathop{\text{Var}}_{L}(S_{t})+\mathop{\text{Var}}_{L}(H_{t})\right]$ (5)
@@
At its core, Dreamer V3 uses a stochastic RSSM to predict the state and reward trajectories over a predefined time horizon given an initial starting state $s_{0}$ and an internal representation $h_{0}$. Due to the stochastic nature of this model, repeating the same trajectory predictions $L\in\mathbb{N}$ times produces a set of state trajectories ($S_{t}=\left\langle s_{t,1},s_{t,2},...,s_{t,L}\right\rangle$) and internal representation trajectories ($H_{t}=\left\langle h_{t,1},h_{t,2},...,h_{t,L}\right\rangle$), each of which provides a plausible estimate of the future states. The variance between trajectories ($\mathop{\text{Var}}_{L}(\cdot)$) thus provides an estimation of the local state divergence as the state perturbation size approaches 0. Therefore, to minimise $\lambda_{1}$ and improve the stability of Dreamer V3 subject to state perturbation, we propose incorporating the regularisation term outlined in Equation 5 into the policy loss (Equation 6). Including this regularisation term as an additional weighted term forces agents to consider the stability of the system during the optimisation process. This incentivises the policy to produce stable state trajectories which attain high rewards instead of solely optimising the expected return. The complete algorithm for calculating the MLE regularisation term is provided in Algorithm 1 using the notation outlined by Hafner et al. [9].
@@ 6 Experiments
In this section, we investigate the impact that the proposed MLE regularisation has on Dreamer V3. We show that the inclusion of this term reduces the chaotic state dynamics produced by the control policy and that this improved stability increases performance when noise is introduced. For these experiments, we train three instances of Dreamer V3 with MLE regularisation and reuse the SAC, TD3 and Dreamer V3 policies from Sections 3 & 4. When estimating state divergence, the RSSM predicts $L=3$ plausible future trajectories over which the state and internal representation variance is measured. Increasing $L$ will produce a more accurate estimate of state divergence; however, this will require more computational resources. Therefore, to maintain a similar training time to that of Dreamer V3, we set $L=3$. All other hyperparameters are consistent with the Dreamer V3 baseline.
@@
Table 2: Average total reward and average MLE produced when Dreamer V3 (DR3) and Dreamer V3 with MLE regularisation (MLE DR3) when controlling various environments sampled from the DeepMind Control Suite [27]. Each policy-environment combination is independently trained with three random seeds using the hyperparameters outlined in Appendix A.2.
@@
Figure 7: Total episode reward for the Cartpole Balance (CB), Cartpole Swingup (CS), Walker Stand (WS), Walker Walk (WW), Walker Run (WR) and Cheetah Run (CR) environments when controlled by trained instances of SAC, TD3, Dreamer V3 (DR3) and Dreamer V3 with MLE regularisation (MLE DR3) subject to $\mathcal{N}(0,\sigma)$ Gaussian observation noise. Each policy-environment combination is independently trained with three random seeds and the average episode reward with a bootstrapped 95% confidence interval is reported over 80 evaluation episodes each with a fixed length of 1000 steps.
To identify how robust these regularised Dreamer V3 policies are to continual state perturbations, we measure their performance when Gaussian noise is added to each observation. Each policy is trained without observation noise ($\sigma=0$) in the fully deterministic variant of each environment and then tested with Gaussian noise ($\mu=0$ and $\sigma\in[0,1]$). Performance is measured over 240 episodes with a fixed length of 1000 steps and a maximum reward per step of 1. Figure 7 shows the interquartile mean and bootstrapped 95% confidence interval for the total reward attained by each policy-environment interaction subject to various levels of Gaussian noise. This shows that MLE regularisation significantly improves the performance of Dreamer V3 in four of the noisy control systems, while the other two environments (Walker Stand and Cheetah Run) attain similar performance to the Dreamer V3 baseline. Furthermore, examining the state trajectories produced by Dreamer V3 and Dreamer V3 + MLE regularisation when controlling the Walker Stand task (Figure 8) shows that the regularisation improves the stability of the control interaction as the trajectories do not diverge significantly. These findings indicate that MLE regularisation can improve the robustness of Dreamer V3 subject to observation noise as it produces more consistent and reliable behaviour in the face of uncertainties.
@@ 7 Conclusion
A key issue preventing the application of deep reinforcement learning to real-world environments is the need for guaranteed stability and performance in the face of noisy observations and adversarial attacks. In this work, we set out to identify the impact a single perturbation has on the long-term behaviour of deep RL policies in continuous control environments. Using the spectrum of Lyapunov Exponents, we established that the MDP can produce chaotic state and reward trajectories which are highly sensitive to initial conditions. This instability poses two threats to the application of deep RL to real-world problems, where it is infeasible to attain an accurate measurement of the system state. First, small state perturbation can have a large impact on the performance of trained deep RL policies, even where the average performance is good. This can create hazards in real-world conditions where observation noise is prevalent and can be exploited by adversarial attack methods. Second, even when deep RL policies perform well, they can produce unpredictable behaviours, which is undesirable in most real-world applications.
To mitigate these chaotic dynamics and improve robustness we propose Maximal Lyapunov Exponent regularisation for Dreamer V3. This novel approach uses the Recurrent State Space Model to estimate the local state divergence and incorporates this into the policy loss. In effect, the agent optimises its confidence in future trajectories jointly with its expectations of rewards. While MLE regularisation helps improve the robustness of the agent’s policies, this approach assumes an accurate estimation of the local state divergence. In environments where the RSSM struggles to capture state dynamics, the effectiveness of the proposed regularisation may be diminished. However, in our experiments, we demonstrate that this regularisation improves the stability of the learnt policies, thereby making them more robust to state perturbations.
@@ Appendix A Appendix
A.1 State space composition
Table 3: State space definition for each environment used in this paper. Each dimension of the state space represents a unique aspect of the control system and can contain any real value. Positions are measured in metres ($m$), angles are measured in radians ($rad$), velocities are measured in metres per second ($m/s$) and angular velocities are measured in radians per second ($rad/s$).
@@ A.1 State space composition
Point mass x & y position
Point mass x & y velocity
@@ B.1 Default values
Table 5: Parameters used to calculate the spectrum of Lyapunov Exponents using the method outlined by Benettin et al. [2, 3].
Parameter Value
@@ B.2 Number of iterations ablation study
Figure 9: Estimated Maximal Lyapunov Exponent of environments sampled from the DeepMind Control Suite when controlled by SAC, TD3 and Dreamer V3 (DR3). Each policy-environment combination is independently trained with three random seeds and evaluated using the parameters in Table 5, except the number of iterations, which varies from 1 to 100. The mean and 95% confidence interval indicate that MLE converges after 100 iterations.
@@ B.3 Number of samples ablation study
Figure 10: Estimated Maximal Lyapunov Exponent of environments sampled from the DeepMind Control Suite when controlled by SAC, TD3 and Dreamer V3 (DR3). Each policy-environment combination is independently trained with three random seeds and evaluated using the parameters in Table 5, except the number of initial samples, which varies from 1 to 20. The mean and 95% confidence interval indicate that MLE converges with 20 initial samples.